OpenStudy (anonymous):

If Sarah is 24 years younger than her mother and if the sum of their ages is 68, ... the expression algbera of thi

OpenStudy (shadowlegendx):

m - 24 = s m + m - 24 = 68 2m - 24 = 68

OpenStudy (shadowlegendx):

The formula was m + s = 68

OpenStudy (shadowlegendx):

So now we have 2m - 24 = 68 Add 24 to both sides Then divide both sides by 2 Tell me what you get